OpenStudy (anonymous):

use summation notation to write the arithmetic series for the specified number of terms then evaluate the sum. 50+55+60+ .....; n=7

OpenStudy (anonymous):

so for this the want you to express it as a single notation \[\sum_{j}^{n}n _{j}\] where n=7

OpenStudy (anonymous):

so you put n = 7 at the bottom right, but how do you do the rest

OpenStudy (anonymous):

summation is \[\sum_{j=50}^{7}(n _{j}+5)\]

OpenStudy (anonymous):

so you start at 50 then add 5 for each subsequent term till you stop at 7 terms

OpenStudy (anonymous):

I gotta go, hopefully this is helpful If not message me and I can look at it again tomorrow

OpenStudy (anikhalder):

\[\sum_{i=0}^{6}(50+5i)\] Using formula : - a=50 d=5 n=7 nth term= \[a+(n-1)d\] (replacing the variables with the actual values) we get nth term(here 7th term) as -> 50+(7-1)5=80 The sum of n terms= \[\frac{ n }{ 2}(2a+(n-1)d)\] (replacing the variables) we get the sum to be \[\frac{ 7 }{ 2}(50\times2 + (7-1)\times5) =455\] Therefore in this A.P., 7th term=80 Summation of the series=455